# A Proxy-Free Strategy for Practically Improving the Poisoning Efficiency in Backdoor Attacks

### v1

@@ Abstract
Poisoning efficiency is a crucial factor in poisoning-based backdoor attacks. Attackers prefer to use as few poisoned samples as possible to achieve the same level of attack strength, in order to remain undetected. Efficient triggers have significantly improved poisoning efficiency, but there is still room for improvement. Recently, selecting efficient samples has shown promise, but it requires a proxy backdoor injection task to find an efficient poisoned sample set, which can lead to performance degradation if the proxy attack settings are different from the actual settings used by the victims. In this paper, we propose a novel Proxy-Free Strategy (PFS) that selects efficient poisoned samples based on individual similarity and set diversity, effectively addressing this issue. We evaluate the proposed strategy on several datasets, triggers, poisoning ratios, architectures, and training hyperparameters. Our experimental results demonstrate that PFS achieves higher backdoor attack strength while 500 faster than previous proxy-based selection approaches.
1 Introduction
@@
The abundance of training data plays a vital role in the success of Deep Neural Networks (DNNs). For instance, GPT-3 [2], a deep learning model with 175 billion parameters, owes its effectiveness in various language processing tasks to its pretraining on a massive corpus of 45 TB of text data. With the increasing demand for data, users and businesses are turning to third-party sources or online collection as easier means of obtaining it. However, recent studies [4, 12, 3, 11] have demonstrated that such practices can be maliciously exploited by attackers to contaminate training data, thereby negatively impacting the functionality and reliability of trained models.
@@
During the training phase, one of the significant threats is called backdoor attacks [4, 12, 27]. This attack involves injecting a covert backdoor into the DNN by introducing a small number of poisoned samples into the benign training set. Although the model appears normal on benign inputs, a predefined trigger can activate the infected model, forcing its predictions to align with the attacker’s objective. As research continues to reveal this threat in various tasks such as speaker verification [46] and malware detection [20], attention to backdoor attacks is increasing.
@@
Poisoning efficiency is a crucial indicator for attackers attempting to launch data poisoning attacks on DNNs [47, 22, 39, 40]. The aim is to contaminate the training dataset with as few poisoned samples as possible while achieving the desired outcome. A perfect scenario is if a backdoor can be successfully implanted by poisoning only one sample in the dataset, significantly enhancing the attack’s stealthiness. Currently, methods to improve poisoning efficiency fall into two categories: designing efficient triggers [47, 48, 44] and selecting efficient samples [39]. The former aims to find trigger patterns that are easier for models to learn, while the latter focuses on selecting and representative poisoned samples. This paper is concerned with the latter approach.
@@
A recent study [39] on sample selection has yielded significant results, demonstrating that the same attack success rate can be achieved with only to of the poisoned samples compared to random selection strategies. However, we have found that this method requires a proxy backdoor injection task to evaluate the contribution of each poisoned sample, resulting in a constructed poisoned set closely related to the attacker’s specific proxy attack. This presents a challenge, as the effectiveness of the selected set will be damaged when the attacker’s proxy attack settings differ from the actual settings of the victims, which is a common scenario in practice. Moreover, conducting the proxy task significantly increases the execution time of the algorithm, limiting its scalability on large datasets.
In this paper, we address the above issue and present a Proxy-Free Strategy (PFS) that acquires higher backdoor attack strength while 500 faster than previous selection method. Our main contributions are as follows:
We conduct an analysis of the forensic features of efficient poisoned samples in backdoor attacks, with a specific focus on two key factors: the similarity between clean and corresponding poisoned samples, and the diversity of the poisoned sample set.
Based on this analysis, we propose a novel selection method called the Proxy-Free Strategy (PFS), which does not require a proxy attack task, thereby addressing the issue of inconsistent settings and significantly reducing the time consumption.
we evaluate PFS on three datasets: CIFAR-10, CIFAR-100, and Tiny-ImageNet, and our results consistently demonstrate its effectiveness. Notably, PFS achieves comparable or even higher attack success rates while 500 faster than previous proxy-based selection approaches, making it more scalable for larger datasets or complex models.
2 Related Works
2.1 Backdoor Attacks
Backdoor attacks aim to embed hidden Trojans into DNNs, enabling the attacked models to perform well on clean samples while misbehaving on samples with specifically designed triggers. These attacks can occur at any stage of DNN development, including code [1, 33], outsourcing, pretrained [36, 10, 19], data collection [12, 4, 35, 26], collaborative learning [42, 31], and post-deployment [32, 17]. Among these, poisoning-based backdoor attacks, which involve injecting backdoor triggers by modifying the training data, are the most straightforward and widely adopted. Recent studies have focused on improving poisoning efficiency, which can be divided into two categories.
@@ 2.1 Backdoor Attacks
Designing Efficient Trigger. Trigger design is a crucial aspect of backdoor attacks. Chen et al. [4] proposed a blended strategy that blends clean samples with the trigger to evade human detection. Later, studies aimed to create more efficient and invisible triggers by adopting natural patterns such as warping [30], rotation [38], style transfer [5], frequency [8, 45], and reflection [28]. In some studies, researchers have taken inspiration from Universal Adversarial Perturbations (UAPs) [29] and optimized an UAP on a pre-trained clean model to create a trigger, which is an effective and widely used approach [48, 21, 6]. Li et al. [23] proposed a different approach where they use GANs models to acquire a sample-specific trigger, which is a departure from previous works that adopt universal triggers.
Selecting Efficient Poisoning Samples. Efficient sample selection for poisoning attacks is a crucial but under-explored area, and is separate from the trigger design. Xia et al. [39] were among the first to explore the contribution of different data to backdoor injection, discovering that not all poisoned samples contribute equally and that appropriate sample selection can significantly improve the efficiency of data in backdoor attacks. However, their study [39] relies on proxy attacks to identify the efficient subset, which suffers from performance degradation when gaps exist (e.g., in Transform (F) and Hyperparameters (H) as shown in Figure 1 between the proxy poisoning attack and the actual poisoning process used by victims. Therefore, a sample selection strategy that is not based on proxy attacks is urgently needed, and is the main contribution of our paper.
2.2 Backdoor Defenses
Backdoor attacks pose a significant threat to the security of DNN models, and several defense methods have been proposed to mitigate these threats. Existing backdoor defenses can be broadly classified into two categories: backdoor detection and backdoor erasing. Backdoor detection aims to detect poisoning in input data or DNN models used [9, 14, 7, 41]. While these methods have demonstrated promising detection precision, they cannot completely mitigate the potential threats of backdoor models. Backdoor erasing aims to purify the threatened models by erasing the backdoor during or post training. During training, clean models can be trained on poisoned data, as proposed by Li et al. [24]. Post-training, models can be fine-tuned, distilled, or pruned to erase the backdoor from the attacked model [25, 37]. These defense methods offer varying levels of effectiveness, and the backdoor defense problem remains an active area of research.
3 Degradation of Proxy-based Selection
@@
This section aims to highlight the limitations of sample selection methods that rely on proxy attacks, motivating the development of a proxy-free approach for selecting poisoned samples in backdoor attacks. To provide context, we will first review the general pipeline of backdoor attacks and the use of proxy attack-based sample selection methods
@@
Poisoning-based Backdoor Attacks. Poisoning-based backdoor attacks involve injecting a backdoor into a subset of a clean training set , resulting in a corresponding poisoned set . Here, is a pre-defined poison generator that adds the trigger into a clean sample , while and represent the true label and attack-target label of the clean sample and the poisoning sample , respectively. After poisoning, the victim will train a deep model:
@@
In the above equation, represents the DNN model, and is the corresponding loss function. The poisoned rate can be calculated as . The goal of sample selection methods is to enhance the efficiency of backdoor attacks by selecting an optimal poisoning set . This approach is orthogonal to the mainstream method of improving poisoning efficiency and has not been thoroughly explored yet.
Sample Selection Based on Proxy Attack. Recent studies [39, 43] have highlighted the varying contributions of different samples to backdoor injection, with certain subsets of efficient data exhibiting higher attack success rates. For instance, [39] suggests that the efficiency of a poisoned samples is related to its forgetting events during backdoor injection. To identify the optimal subset, they formulate sample selection as an optimization problem and propose a Filtering-and-Updating Strategy (FUS). Although FUS achieves remarkable performance, it utilizes a proxy attack processing to identify the efficient subset.
@@
As illustrated in Figure 1, poisoning-based backdoor attacks consist of two phases: The attacker combines data (D) and trigger (T) to generate poisoned data (D+T) and releases it to victims, who then use it to train infected DNN models. During training, various data transforms/augmentations (F) and hyperparameters (H)***In this case, H includes any training settings such as the model architecture, the optimizer, and the training hyperparameters. are applied. Typically, the attacker has no access to any information about the victim phase. However, proxy-based sample selection methods rely on the complete attack process to select samples. We argue that all components involved in backdoor attacks, including D, T, F, and H, can affect the efficiency of poisoned samples. When discrepancies arise (e.g., between the transforms and hyperparameters used by the attacker and those used by victims), the effectiveness of the selected samples may be compromised. Therefore, we will first investigate the impact of such discrepancies on proxy-based sample selection methods.
3.1 Experimental Settings
@@
To investigate the proposed question in this section, we conduct experiments using the VGG-16 model [34] as the victim model on the CIFAR-10 dataset [18]. We use the SGD optimizer with an initial learning rate of 0.01, momentum of 0.9, and weight decay of 5e-4, and apply two common data transformations: random crop and random horizontal flip. The training is performed for 70 epochs with a batch size of 256. Poisoned samples are generated using a blending strategy [4] in which the clean samples are blended with a trigger image using the equation , where is set to 0.15. We set the attack target to category 0. We adopt the FUS method with a poisoned rate (the size of the poisoning set is ) and the number of selection iterations following [39]. In accordance with previous literature [22], the attack success rate is defined as the probability of classifying a poisoned test data to the target label. Unless stated otherwise, all experiments in this section follow the aforementioned settings.
3.2 Empirical Studies
@@
Different Data Transforms (F) will decrease the effectiveness of proxy-based sample selection method. We investigated how different data transforms (F) affect the effectiveness of proxy-based sample selection methods. To conduct our analysis, we used the experimental settings described previously and followed them exactly. Specifically, we set the data transform for the proxy poisoning attack in the sample selection phase (FUS) to ”RandomCrop” and ”RandomHorizontalFlip”, and we used ”None”, ”RandomCrop”, ”RandomHorizontalFlip”, ”RandomRotation”, and ”ColorJitter” for the data transform during the actual poisoning attack in the victim phase. Our results, depicted in Figure 2, indicate that different data transforms (F) decrease the effectiveness of proxy-based sample selection methods.
Different Hyperparameter (H) will decrease the effectiveness of proxy-based sample selection method. Furthermore, we explore the impact of different hyperparameter to efficient data searched by FUS [39]. Following the same experimental settings as described above, we conduct experiments using the hyperparameter ”V-16, SGD, 0.01” to search for efficient sets of backdoor attacks and evaluate the attack success rate on several other hyperparameters, including ”V-13, SGD, 0.01”, ”V-13, Adam, 0.001”, ”V-16, Adam, 0.001”, ”R-18, SGD, 0.01”, ”R-18, Adam, 0.001”, and ”P-18, SGD, 0.01”†††V-13, V-16, R-18, and P-18 denote VGG13, VGG-16, ResNet-18, and PreActResNet-18, respectively. . Our results, shown in Figure 3, demonstrate that different hyperparameters (H) can also significantly decrease the effectiveness of the proxy-based sample selection method.
Overall, our results suggest that the efficacy of proxy-based sample selection methods is sensitive to attacker-agnostic data transforms and hyperparameters used during the victim phase. This highlights the importance of using proxy-free search methods to achieve efficient sample selection in backdoor attacks.

### v2

# A Proxy Attack-Free Strategy for Practically Improving the Poisoning Efficiency in Backdoor Attacks

@@ Abstract
Poisoning efficiency plays a critical role in poisoning-based backdoor attacks. To evade detection, attackers aim to use the fewest poisoning samples while achieving the desired attack strength. Although efficient triggers have significantly improved poisoning efficiency, there is still room for further enhancement. Recently, selecting efficient samples has shown promise, but it often requires a proxy backdoor injection task to identify an efficient poisoning sample set. However, the proxy attack-based approach can lead to performance degradation if the proxy attack settings differ from those used by the actual victims due to the shortcut of backdoor learning. This paper presents a Proxy attack-Free Strategy (PFS) designed to identify efficient poisoning samples based on individual similarity and ensemble diversity, effectively addressing the mentioned concern. The proposed PFS is motivated by the observation that selecting the to-be-poisoned samples with high similarity between clean samples and their corresponding poisoning samples results in significantly higher attack success rates compared to using samples with low similarity. Furthermore, theoretical analyses for this phenomenon are provided based on the theory of active learning and neural tangent kernel. We comprehensively evaluate the proposed strategy across various datasets, triggers, poisoning rates, architectures, and training hyperparameters. Our experimental results demonstrate that PFS enhances backdoor attack efficiency, while also exhibiting a remarkable speed advantage over prior proxy-dependent selection methodologies.
Index Terms:
Backdoor attack, Sample selection.I Introduction
@@
The abundance of training data plays a vital role in the success of Deep Neural Networks (DNNs). For instance, GPT-3 [3], a deep learning model with 175 billion parameters, owes its effectiveness in various language processing tasks to its pretraining on a massive corpus of 45 TB of text data. Responding to the increasing demand for data, both users and businesses are increasingly turning to third-party sources or online repositories, which offer more convenient way for data collection. However, recent researches [4, 14, 13] have demonstrated that such practices can be maliciously exploited by attackers to poison training data, thereby negatively impacting the functionality and reliability of trained models.
@@
During the training phase, a significant threat emerges in the form of backdoor attacks [4, 14, 33]. This attack involves injecting a covert backdoor into the DNNs by introducing a small number of poisoning samples into an otherwise benign training dataset. While the model appears normal when presented with benign inputs, a predefined trigger can activate the infected model, forcing its predictions to align with the objective of attackers. As research continues to reveal this threat in various tasks such as speaker verification [59] and malware detection [24], attention to backdoor attacks is increasing.
@@
Poisoning efficiency stands as a pivotal metric for attackers engaging in backdoor attacks against DNNs [60, 26, 50, 52]. Their objective revolves around poisoning the training dataset with minimal poisoning samples while achieving the desired outcomes. An ideal scenario is one where a single poisoning sample suffices to implant a backdoor, greatly enhancing the stealthiness of the attack. Currently, strategies aimed at enhancing poisoning efficiency fall into two main categories: designing efficient triggers [60, 61, 57] and selecting efficient samples [50]. The former focuses on identifying trigger patterns that are easier for models to learn, while the latter centers around the careful selection of representative samples for poisoning. This paper aligns with the latter approach, exploring the differences in poisoning efficiency of different poisoning samples in backdoor attacks.
@@
A recent investigation [50] into sample selection has yielded significant results, demonstrating that an equivalent attack success rate can be achieved using only to of the poisoning samples in comparison to random selection strategies. However, it has become apparent that this method relies on a proxy backdoor task to assess the contribution of each poisoning sample, thus constructing an efficient poison set closely associated with the specific proxy attack of the attacker. This proxy attack-based selection method [50] poses a challenge due to potential deterioration in its effectiveness when the proxy attack settings of attackers deviate from the actual settings of the victims—a frequent attack scenario [60, 61, 57]. Moreover, conducting the proxy attack task significantly increases the execution time of the algorithm, limiting its scalability on large datasets.
This paper presents a solution to the aforementioned challenge by introducing a novel Proxy attack-Free Strategy (PFS) that not only achieves higher backdoor attack efficiency but also boasts a remarkable speedup compared to previous selection methods. Our contributions can be summarized as follows:
This paper makes a comprehensive study of the forensic properties inherent in efficient poisoning samples within the context of backdoor attacks. In particular, we delve into two critical facets: the similarity between clean samples and their corresponding poisoning samples, and the diversity of poisoning set.
Drawing on the insights derived from our analysis, we introduce a novel selection approach named the Proxy attack-Free Strategy (PFS). Distinguishing itself from previous study, PFS eliminates the requirement for a proxy attack task, effectively addressing the challenge arising from differing settings while significantly reducing the time overhead.
The effectiveness of PFS is subjected to evaluation across three diverse datasets: CIFAR-10, CIFAR-100, and Tiny-ImageNet. Our empirical findings consistently confirm its efficacy. It is worth noting that PFS has achieved superior attack success rates while achieving a remarkable acceleration when contrasted with prior proxy-dependent selection methodologies. This attribute renders PFS particularly well-suited for scalability, making it apt for accommodating larger datasets and complex models.
II Related Works
II-A Backdoor Attacks
Backdoor attacks aim to inject Trojans within DNNs. This manipulation empowers poisoning models to yield accurate outcomes when presented with clean samples, while displaying anomalous behavior with samples containing designed triggers. These attacks can occur at various stages within the DNN development lifecycle, including code [1, 41], outsourcing [51], pre-trained models [45, 12], data collection [14, 4, 44, 31], collaborative learning [54, 39], and even post-deployment scenarios [40, 21]. Of these ways, the most straightforward and widely adopted approach is poisoning-based backdoor attacks, entailing the insertion of backdoor triggers through modifications to the training data. Recent researches on poisoning-based backdoor attacks have been proposed to enhancing poisoning efficiency from two aspects.
@@ II-A Backdoor Attacks
Designing Efficient Triggers. Trigger design is a hot topic in backdoor learning. Chen et al. [4] initially proposed a blended strategy for evading human detection, acquiring poisoning samples by blending clean samples with triggers. Subsequent studies have explored leveraging natural patterns such as warping [38], rotation [48], style transfer [5], frequency [9, 58], and reflection [34] to create triggers that are more imperceptible and efficient compared to previous methods. Drawing inspiration from Universal Adversarial Perturbations (UAPs) [37] in adversarial examples, some research [61, 25, 7] optimizes an UAP on a pre-trained clean model as the trigger, a highly effective and widely adopted approach. Unlike previous approaches that use universal triggers, Li et al. [27] employ GANs models to generate sample-specific triggers.
Selecting Efficient Poisoning Samples. Efficient poisoning samples selection is an under-explored area and is orthogonal to the trigger design. Xia et al. [50] first explored the contribution to backdoor injection for different data. Their work illustrated the inequitable influence of individual poisoning samples and underscored the enormous potential for improved data efficiency within backdoor attacks through efficient sample selection. Simultaneously, Gao et al. [10] assert a similar notion—that not all samples equally foster the process of poisoning in backdoor attacks. They introduce the concept of ”forgetting events” to spotlight the most potent poisoning samples in the context of clean-label backdoor attacks. However, both of these studies [50, 10] rely on proxy attacks to determine the efficient subset of samples. Regrettably, this dependence on proxy attacks introduces a vulnerability, leading to a degradation in attack performance when disparities arise in parameters like Transformation (F) and Hyperparameters (H) between the proxy poisoning attack and the actual poisoning process, as depicted in Figure 1. Therefore, a sample selection strategy that is not based on proxy attacks is urgently needed, and is the main contribution of our paper. Comparable to our approach, the Outlier Poisoning Strategy (OPS) [17] also utilizes a surrogate model for efficient sample selection in backdoor attacks against anti-spoof rebroadcast detection. The primary concept behind this strategy involves embedding triggers into samples that pose the greatest classification challenge (referred to as outlier samples), encouraging the network to depend on the trigger for classification. However, OPS [17] is specifically designed for binary classification and clean-label settings, exhibiting diminished performance in scenarios involving dirty-label settings and multiple classifications.
II-B Backdoor Defenses
Backdoor attacks pose a significant threat to the security of Deep Neural Network (DNN), leading to the emergence of various defense mechanisms aimed at mitigating these risks. The current landscape of backdoor defenses can be broadly categorized into two classes: backdoor detection and backdoor erasure. Backdoor detection aims to identify the presence of poisoning either in input data or within the employed DNN models themselves [11, 18, 8, 53]. Conversely, backdoor erasure strategies aim to repair poisoning models by removing the embedded backdoor, either during or after the training process. In the training phase, the concept of training clean models on poisoned data has been introduced, exemplified by Li et al. [28]. Post-training, a range of techniques, including fine-tuning, distillation, and pruning, can be employed to eliminate the backdoor from the poisoning model [29, 47].
III Degradation of proxy attack-based Selection
@@
The purpose of this section is to emphasize the limitations associated with sample selection methods that rely on proxy attacks, which serves as a motivation for designing a proxy attack-free sample selection strategy in backdoor attacks. To better understand our key observations, we will first review the pipeline of backdoor attacks and sample selection methods based on proxy attack.
@@
Poisoning-based Backdoor Attacks. Poisoning-based backdoor attacks involve injecting a backdoor into a subset of a clean training set , resulting in a corresponding poisoned set . Here, is a pre-defined generator that adds the trigger into clean sample , while and represent the true label and attack-target label of the clean sample and the poisoning sample , respectively. After poisoning, the victim will train a deep model:
@@
In the above equation, represents the DNN model, and is the corresponding loss function. The poisoning rate can be calculated as . The goal of sample selection methods is to enhance the efficiency of backdoor attacks by selecting an optimal poisoning set . This approach is orthogonal to the mainstream method of improving poisoning efficiency and has not been thoroughly explored yet.
Proxy Attacks-based Sample Selection. Recent investigations [50, 55, 10] have highlighted that diverse samples contribute differently to the backdoor injection and a particular subset with efficient data has higher attack success rate. Notably, [50, 10] claim that the efficacy of poisoning samples is intertwined with instances of ”forgetting events” during the backdoor injection process. To distill the most efficient subset, [50]***It is worth noting that both [50] and [10] employ the concept of forgetting events to guide the selection of poisoned sets. It is worth highlighting that while FUS [50] employs multiple iterations of screening, [10] adopts a one-step screening approach, rendering [10] a specific case of FUS. As such, for the purpose of analyzing the drawbacks of proxy attack-based methods for selecting poisoning samples, we treat FUS as a representative example in this section. formulates the sample selection as an optimization problem and introduce a Filtering-and-Updating Strategy (FUS). Although FUS has acquired remarkable performance, it uses a proxy attack processing to find the efficient subset.
@@
As illustrated in Figure 1, poisoning-based backdoor attacks contain two phases: The attacker combines data (D) and trigger (T) to build poisoned data (D+T) and release it. Victims download the released data and adopt it to train infected DNN models. Notably, during the training process, various data transformations/augmentations (F) and hyperparameters (H)†††In this context, H encompasses various training settings, encompassing elements like the model architecture, optimizer choice, and training hyperparameters. have been applied by victims. Generally, the attacker has no access to any information of the victim phase and only performs the poisoning injection at the attack phase. However, proxy attack-based sample selection methods need to select samples with the help of the complete attack process (both attacker and victim phases). We argue that all parts involved in backdoor attacks are the factors that affect the efficiency of poisoning samples, containing D, T, F, and H. When there contains some gaps (e.g. Transform (F) and Hyperparameters (H) in Figure 1) between proxy poisoning attack and actual poisoning process of victims, the poisoning effectiveness of the selected samples may be decreased. Following, we will explore what will happen if the actual poisoning process is different from the proxy poisoning attack in proxy attack-based sample selection methods.
III-A Experimental Settings
@@
To investigate the proposed question within this section, we engage in a series of experiments utilizing the VGG16 model [43] as the victim model. These experiments are conducted on the CIFAR-10 dataset [23]. Our chosen config encompasses the SGD optimizer, initially set with a learning rate of 0.01, a momentum value of 0.9, and a weight decay coefficient of 5e-4. Notably, we incorporate two widely adopted data transformations—random cropping and random horizontal flipping. The training is performed for 70 epochs with a batch size of 256. Poisoning samples are generated using a blending strategy [4] in which the clean samples are blended with a trigger image using the equation , where is set to 0.15. We set category 0 as the target for the attack, and we set the poisoning rate with —equating to a poisoning set size of . The process involves 15 iterations of selection, aligning with the settings of [50]. In accordance with previous literature [26], the attack success rate is defined as the probability of classifying a poisoned test data to the target label. Unless stated otherwise, all experiments in this section follow the aforementioned settings. To explore the degeneration in poisoning efficiency within selected samples due to disparities between the proxy poisoning attack and the actual poisoning process of victims, we compare the Filtering-and-Updating Strategy (FUS) with SELF-FUS. SELF-FUS represents an ideal scenario for FUS, assuming the attacker possesses complete information about both the attacker and victim phases during poisoning-based backdoor attacks—a scenario that doesn’t align with the threat model in this study. Therefore, proxy poisoning attack of attacker phase is consistent with the actual poisoning process of victim phase in SELF-FUS, which is not guaranteed in the FUS.
III-B Empirical Studies
@@
Impact of Different Data Transformations (F) on Proxy Attack-based Sample Selection Efficacy. We investigated how different data transforms (F) affect the effectiveness of proxy attack-based sample selection methods. To conduct our analysis, we used the experimental settings described previously and followed them exactly. Specifically, we configured the data transformations for the proxy poisoning attack during the sample selection phase (FUS) as ”RandomCrop” and ”RandomHorizontalFlip”. Subsequently, during the actual poisoning attack within the victim phase, we introduced variations by employing a range of transformations, including ”None”, ”RandomCrop”, ”RandomHorizontalFlip”, ”RandomRotation”, and ”ColorJitter”. Our results, depicted in Figure 2, indicate that different data transforms (F) decrease the effectiveness of proxy attack-based sample selection methods. An exception to this pattern emerges in cases where the settings for both the proxy and actual attacks align closely, exemplified by the ”1+2” searching and ”1” attacking scenario. This particular outcome is logical and can be attributed to the inherent margin of error.
Effect of Different Hyperparameters (H) on Proxy Attack-based Sample Selection Efficiency. In this section, We thoroughly investigated the impact of different hyperparameters on the effective data efficiency searched by FUS [50]. Adhering to the aforementioned experimental framework, we configured the hyperparameter combination for the proxy poisoning attack during the sample selection phase as ”V-16, SGD, 0.01”. Subsequently, during the actual poisoning attack within the victim phase, we introduced variations by employing a range of hyperparameters, including ”V-13, SGD, 0.01”, ”V-13, Adam, 0.001”, ”V-16, Adam, 0.001”, ”R-18, SGD, 0.01”, ”R-18, Adam, 0.001”, and ”P-18, SGD, 0.01”‡‡‡In this context, V-13, V-16, R-18, and P-18 denote VGG13, VGG16, ResNet18, and PreActResNet18, respectively. . Figure 3 indicates that different hyperparameters (H) decrease the effectiveness of proxy attack-based sample selection methods. Once again, we observed anomalous occurrences when the parameters for both proxy searching and actual attacks closely converged, as illustrated by the scenario involving ”V-16-SGD-0.01” proxy searching and subsequent ”V-13-SGD-0.01” attacks within the victim phase.
Summary Collectively, our findings underscore the susceptibility of proxy attack-based sample selection techniques to shifts in data transforms and hyperparameters deployed during the victim phase. Notably, SELF-FUS consistently outperforms FUS across various scenarios, with the exception of outliers observed in both Figure 2 and Figure 3. These outliers, marked by the ”1+2” searching and ”1” attacking in Figure 2, as well as the ”V-16-SGD-0.01” searching and ”V-13-SGD-0.01” attacking in Figure 3, share closely aligned proxy and actual attack settings. To further validate our hypothesis, we continue to simulate more realistic attack scenarios that introduce disparities in both data transformations (F) and hyperparameters (H) between proxy and actual attacks. Figure 4 shows the ASR comparison between FUS and SELF-FUS on both CIFAR-10 and Tiny-ImageNet datasets. The notable difference in ASR between FUS and SELF-FUS underscores the significant impact of proxy attack-based methods on sample efficiency. Specifically, it highlights how FUS experiences serious degradation in sample efficiency when there is a substantial difference between the settings for proxy poisoning attacks during the attacker phase and those employed in the actual poisoning process during the victim phase.
IV Methodology
To improve poisoning efficiency, it is important to develop proxy attack-free sample selection methods that leverage the forensic features of efficient data in poisoning-based backdoor attacks. Before introducing our method, we will first discuss some key observations.
IV-A Forensic Features of Efficient Data in Poisoning-based Backdoor Attacks
We assert that the primary reason for the efficiency of backdoor injection is due to the similarity between benign and corresponding poisoning samples. Typically, poisoning samples exhibit different labels than their clean counterparts, signifying that these pairs share similar attributes yet are categorized differently in dirty-label backdoor attacks. In this context, samples exhibiting a high degree of similarity can be considered challenging samples in the context of poisoning tasks. Consequently, they tend to be more efficient than low-similarity samples for the purpose of backdoor attacks. To substantiate this assumption, we conducted empirical investigations utilizing the CIFAR-10 dataset while adhering to the same experimental settings as detailed in Sec. III-A. The left segment of Figure 5 visually captures the distribution of similarities among 500 randomly chosen samples and 500 samples searched using the FUS method [50]. The result demonstrates that the FUS-searched efficient sample exhibits higher similarity compared to randomly selected samples.
We proceeded to explore the contribution of similarity between benign and corresponding poisoning samples to the efficacy of backdoor attacks. The right part of Figure 5 showcases the attack success rate across varying sets of poisoning samples. Notably, our observations align with our expectations: using high-similarity samples (Top-500 similarity sampling) for poisoning yielded significantly higher attack success rates than utilizing low-similarity samples (Bottom-500 similarity sampling). Furthermore, we compared random sampling with similarity-based selection. However, both high-similarity samples (Top-500 similarity sampling) and low-similarity samples (Bottom-500 similarity sampling) for poisoning yielded lower attack success rates compared to random sampling (Random sampling). This implies that while similarity can effectively filter out inefficient samples, it lacks a positive correlation with data efficiency in the context of poisoning attacks.
We argue that the degeneration observed in the performance of the high-similarity samples (Top-500 similarity sampling) can be attributed to the excessively constrained diversity within the pool of poisoning samples. The middle part of Figure 5 further illustrates this point by displaying the t-SNE representation of the entire 50,000-sample dataset and the 500 selected samples. It is evident that the high-similarity samples exhibit limited diversity. To address this limitation, we sampled 500 specimens randomly from the top 5000 samples with the highest similarity (Top-5K similarity random 500 sampling). This alleviates the diversity constraint, as confirmed by the right segment of Figure 5. Among these approaches, the Top-5K similarity random 500 sampling method is more efficient than random sampling method.
In summary, the similarity between benign and corresponding poisoning samples and the diversity within the poisoning sample set emerge as two pivotal factors influencing the sample efficiency in backdoor attacks.
IV-B Threat Model
Attacker’s Capacities. In accordance with the fundamental prerequisites of poisoning-based backdoor attacks [26, 50], we assume that attackers possess the ability to poison a fraction of the training dataset. However, they do not possess access to any additional training components encompassed within the victim phase, including the training loss, schedule, or model architecture. This assumption describes a more challenging and realistic scenario [26, 46], illustrating the attackers’ constrained knowledge about the target system. Furthermore, we also assume that attackers require a pre-trained feature extractor derived from the accessed training dataset. This assumption is widely acknowledged within the current backdoor attacks [61, 26].
Attacker’s Goals. Our objective is to augment the effectiveness of dirty-label poisoning backdoor attacks costless, achieved through the judicious selection of a suitable poisoning set from the dataset . This approach is orthogonal to most of the current methods for improving poisoning efficiency through designing trigger. Therefore, our method can be integrated to other attacks at a fraction of the cost.